Assume that the risk adjusted cost of capital is 16% for project A, 18% for project...

  1. Assume that the risk adjusted cost of capital is 16% for project A, 18% for project B, and 20% for project C.


Project A:              Estimated cash flow is $25,800 per year for 10 years.
                              Cost $100,000


Project B:              Estimated cash flow is $33,500 per year for 8 years.
                              Cost $125,000


Project C:              Estimated cash flow is $44,000 per year for 7 years.
                              Cost $155,000


A. Compute the Payback, NPV, IRR and Profitability Index for each.
B. Rank order the Projects according to the value to the firm
Answer #1

A

Project A
Year Cash flow stream Cumulative cash flow
0 -100000 -100000
1 25800 -74200
2 25800 -48400
3 25800 -22600
4 25800 3200
5 25800 29000
6 25800 54800
7 25800 80600
8 25800 106400
9 25800 132200
10 25800 158000
Payback period is the time by which undiscounted cashflow cover the intial investment outlay
this is happening between year 3 and 4
therefore by interpolation payback period = 3 + (0-(-22600))/(3200-(-22600))
3.88 Years
Project A
Discount rate 0.16
Year 0 1 2 3 4 5 6 7 8 9 10
Cash flow stream -100000 25800 25800 25800 25800 25800 25800 25800 25800 25800 25800
Discounting factor 1 1.16 1.3456 1.560896 1.8106394 2.100342 2.436396 2.82622 3.278415 3.802961 4.411435
Discounted cash flows project -100000 22241.38 19173.6 16528.97 14249.11 12283.72 10589.41 9128.802 7869.657 6784.187 5848.437
NPV = Sum of discounted cash flows
NPV Project A = 24697.27
Where
Discounting factor = (1 + discount rate)^(Corresponding period in years)
Discounted Cashflow= Cash flow stream/discounting factor
Project A
IRR is the rate at which NPV =0
IRR 22%
Year 0 1 2 3 4 5 6 7 8 9 10
Cash flow stream -100000 25800 25800 25800 25800 25800 25800 25800 25800 25800 25800
Discounting factor 1 1.223746 1.497555 1.832628 2.2426724 2.744462 3.358526 4.109985 5.029579 6.15493 7.532074
Discounted cash flows project -100000 21082.8 17228.08 14078.14 11504.132 9400.748 7681.941 6277.396 5129.654 4191.762 3425.351
NPV = Sum of discounted cash flows
NPV Project A = 7.92529E-05
Where
Discounting factor = (1 + IRR)^(Corresponding period in years)
Discounted Cashflow= Cash flow stream/discounting factor
IRR= 22.37%
Project A
PI= (NPV+initial inv.)/initial inv.
=(24697.27+100000)/100000
1.25



B

Based on NPV

Rank 1 Project A

Rank 2 Project B

Rank 3 Project C